OpenStudy (anonymous):

Suppose Y~p(y)=1/30(y-5)^2 for y = 1,2,3,4. Calculate E(1/Y).

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

I don't even know how to start.

OpenStudy (zarkon):

do you know how to find expected values?

OpenStudy (anonymous):

the summation of the y*p(y)

OpenStudy (zarkon):

\[E(G(Y))=\sum_{y\in\Omega}G(y)P(Y=y)\]

OpenStudy (zarkon):

here your \(G(y)=\frac{1}{y}\)

OpenStudy (anonymous):

so the summation should be: \[\sum_{?}^{?}\frac{ 1 }{ y }(\frac{ 1 }{ 30 })(y-5)^2 \] ?

OpenStudy (zarkon):

\[\sum_{y=1}^{4}\frac{ 1 }{ y }\left(\frac{ 1 }{ 30 }\right)(y-5)^2\]

OpenStudy (anonymous):

right, okay. my prof also gave me r=1-p. Is that going to be useful at all? or is it just extra information?

OpenStudy (zarkon):

I would just compute the above

OpenStudy (anonymous):

okay. Thank you.

OpenStudy (zarkon):

I get 53/72

OpenStudy (anonymous):

I got the same. Thank you again!